Chirag’s Chapo Challenge Chirag, one of the senior-most members of our club, has recently started getting his hands dirty with the markets. He decided to invest the prize money he got by winning the Auquan Quant challenge into a trading strategy. He invested INR 10,000 that would give him returns of 10% per annum. (strategy compounds money on a daily basis) You, a young chap, highly motivated to pursue finance, went to him seeking guidance, so as to how to start learning Finance. Along with a gyaan session, comes an unsaid CHAPO, right! But, but CHAPO doesn’t come so easily, facche!. So, Chirag gives you the money he wanted to invest in the market, and proposes you a deal! You get a Chapo after the recruitment process of the club if you could make more money than the strategy he wanted to invest in. If you are successful in making more money than that strategy, the invested capital is all yours (i.e., Chapo of INR 10,000) and Chirag takes the return (money you made from the strategy) home. But, but. If your strategy gives return less than 10%, you return the money to Chirag, along with the return that you generated from your strategy and also owe the amount that you were not able to make. (Say if you created 4% return, you’d have to give INR 10400 plus the remaining 6% to Chirag from your pocket, and you don’t get a Chapo, of course!) Assuming that you have been following the news for a while, you recall that banks in Argentina offer a 20% interest rate per annum (given: the bank also compounds money on a daily basis) on the money (in Peso, of course) deposited. (Woah, a bank giving more return than a sophisticated trading strategy used by Chirag!). So with this knowledge, you decided to keep that money in an Argentinian bank. Impressed by your idea of investing in the foreign market, Chirag added on: “Take this information from me, you’ll need this. FX markets aren’t so easy as they look.” Spot rate: 1.20 INR/ARS Forward rate (spot rate after one year): 1.05 INR/ARS He suggested “If you were to take my suggestion, keep the money in the bank for a year, don’t withdraw it before one year” Chirag is witty and smart, inherently. The Chapo won’t come so easy! This time he would have made some errors of judgment, and you could get a grand CHAPO. Maybe, or may not. Should you follow his suggestion? Justify your choice. Suppose if you follow his suggestion, and get into the contract, calculate the money that Chirag would get/ or the money which you have to pay from the pocket, depending upon the returns from the contract. Justify your answer. No matter how witty he is, we are sure he wouldn’t propose a deal that would make taking a Chapo impossible for you, we being confident about how kind-hearted he is! So, you have the flexibility to invest the money in the Argentinian bank for the period you want (but not more than a year). The spot rate of INR/ARS follows this function: Spot rate = 1.20 (t=0) Spot rate = 2e-0.01t (0
Answer #1

The suggestion of keeping the money in the Argentinian bank for a year should not be followed because of the following

INR 10,000 converted to ARS gives 10000/1.2 = ARS 8333.33

This money invested in the Argentinian Bank @20% after one year becomes

=8333.33 * (1+0.2/365)^365 = ARS 10177.80   

After one year, the above amount in Argentinian Currency can be converted back to INR at the rate of iNR1.05/ARS

So,   after one year , you get INR 10177.80 *1.05 = INR 10686.69

This gives an effective return of around 6.8%  only on the INR ,, So the CHAPO is not possible by this strategy.

By Chirag's strategy. the amount would be = 10000* (1+0.1/365)^365 after one year =INR 11051.56

So, the shortfall would be 11051.56 - 10686.69 = INR 364.87

So, you would have to pay all the money along with INR 364.87 from your own pocket.

** Please note that with daily compounding, the above answer is correct. However, one can closely approximate the same with continuous compounding formula also.